OpenStudy (anonymous):

what is an equation for the line that passes through points (8, -3) and (9, 10)

OpenStudy (anonymous):

Ok hold let me see wasssup do you mutiple choice?

OpenStudy (anonymous):

have

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

a. (y+3)=-13 (x-8) b.(y+3)=13(x-8) c.(y-3)=-13(x=8) d.(y-3)=13(x+8)

OpenStudy (anonymous):

Ok hold on give me 2 minutes okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I tried my best ion know if it's right but, the answer I got was B

OpenStudy (anonymous):

was that the answer?

OpenStudy (anonymous):

ill tell u in like 3 minutes okay

OpenStudy (anonymous):

okaay